# Functional Stochastic Gradient MCMC for Bayesian Neural Networks

### v1

@@ 5.2 UCI Regression
In this experiment, we evaluate our method for multivariate regression tasks on 7 benchmark UCI datasets: Yacht, Boston, Concrete, Energy, Wine, Kin8nm and Protein. We use a 2- hidden-layer fully connected neural network. For all sampling methods, we run 500 iterations for the burn-in stage and collect 15 samples. Performance is measured via average root mean square error (RMSE) and the test negative log-likelihood (NLL). For each dataset, we construct 10 random 90-to-10 train-test splits and report the mean and standard deviation of RMSE and NLL over the splits in Table 2 and Table 3, respectively. fSGHMC outperforms all baselines for RMSE on all datasets, and fSGLD also achieves better accuracy values than naive SGMCMC methods and functional variational inference, demonstrating superior performance of our functional SGMCMC in predictive accuracy. Moreover, fSGHMC and fSGLD obtain better NLL results on most of the datasets (5/7), showing that our method is also very competitive in uncertainty quantification.
Table 2: The table shows the results of average RMSE for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
@@
Table 3: The table shows the results of average NLL for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
@@ 5.3 Imagine Classification and OOD Detection
In this section, we investigate the scalability of functional SGMCMC on image classification tasks with high-dimensional inputs. We evaluate the in-distribution predictive performance and out-of-distribution (OOD) detection ability on MNIST (LeCun et al., 2010), FashionMNIST(Xiao et al., 2017) and CIFAR-10(Krizhevsky et al., 2009). We compare our fSGLD and fSGHMC with naive SGLD, SGHMC, functional variational FBNN and IFBNN. We use ResNet-18 architecture (He et al., 2016) for all methods on these three datasets, and the batch size is 125. For all sampling methods, we run for 60 burn-in iterations and collect 10 samples on the MNIST dataset, 100 burn-in iterations and collect 10 samples on FashionMNIST, 80 burn-in iterations and draw 10 samples on CIFAR10. The functional GP prior with the RBF kernel is pre-trained on the test dataset for all functional methods. We randomly sample 40 points from training data in every iteration as the measurement points for the gradient estimation of log functional prior distribution in fSGLD and fSGHMC. We report the test classification errors for predictive performance and the area under the curve (AUC) of OOD detection pairs FashionMNIST/MNIST, MNIST/FashionMNIST and CIFAR10/SVNH based on predictive entropies in Table 4. Our fSGLD and fSGHMC outperform all other methods for both predictive accuracy and OOD detection.
@@ 5.4 Contextual Bandits
In many downstream tasks, such as online sequential decision-making scenarios in contextual bandits problems, reliable uncertainty estimation is important to guide the exploration-exploitation process. In such problems, an agent interacts with an unknown environment repeatedly and chooses to take an optimal action in each round of interaction. Thompson sampling (Thompson, 1933; Russo & Van Roy, 2016) is a widely used algorithm for exploration strategies in contextual bandits, it (i) first samples a current posterior to obtain a model configuration; (ii) then adaptively chooses an optimal action based on the current context under the sampled model configuration and observes the corresponding reward; (iii) updates the posterior based on the context, action and reward tuples in this interaction round.
This section evaluates the ability to guide exploration on the UCI Mushroom dataset with 8124 instances. In each instance, the mushroom has 22 features and is labeled as edible or poisonous. An agent can observe mushroom features as the context in each interaction and choose to eat or reject a mushroom. We compare our fSGLD with several baselines, including naive SGLD, functional variational FBNN and IFBNN, parametric-space variational method Bayes By Backprop (BBB) (Blundell et al., 2015). We follow the basic settings by Blundell et al. (2015) and consider three different reward patterns: if the agent eats an edible mushroom, it receives a reward of 5 and a reward of 0 if the agent chooses to reject. On the other hand, if the agent eats a poisonous mushroom, it receives a reward of -35 with probabilities of 0.4, 0.5, and 0.6, respectively, for three different patterns. If the agent chooses to reject a poisonous mushroom, it receives a reward of 0. Suppose an oracle always chooses to eat an edible mushroom and reject the poisonous mushroom. The cumulative regrets with respect to the reward achieved by the oracle measure the exploration-exploitation ability of an agent. We run 10000 iterations for all methods and the minibatch size is 64. Figure 2 shows the cumulative regrets for all methods in three reward patterns. Our fSGLD consistently outperforms other inference methods, indicating that our functional SGMCMC can provide reliable uncertainty estimation in decision-making tasks.
@@ 6 Discussion and Conclusions
In this paper, we have lifted the Itô diffusion of network weights in BNNs to function space and proposed a new functional SGMCMC scheme for BNNs. We have designed the functional Langevin dynamics and Hamiltonian dynamics for the function mapping defined by BNNs directly, which can incorporate informative functional prior into posterior inference. Moreover, we proved that the stationary distribution of our functional SGMCMC is the posterior distribution over functions. Empirically, we demonstrated that our functional SGMCMC approaches can leverage functional priors to yield high predictive performance and principled uncertainty quantification on several benchmark tasks.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Further Background
@@
Itô Lemma (Itô, 1951) is a fundamental result in stochastic calculus to find the differential of a function of a stochastic process, which serves as the stochastic calculus counterpart of the chain rule. For an Itô diffusion, let $f(X_{t})$ be an arbitrary twice differentiable scalar function of real variables $X_{t}$, then the differential of $f(X_{t})$ can be derived from the Taylor series expansion of the function as
@@
which implies that $f(X_{t})$ is itself an Itô diffusion (Brzeźniak et al., 2008).
Appendix B Theoretical Proof
The Fokker–Planck equation (Risken, 1996) for the corresponding evolution of the probability density $p(x,t)$ of a diffusion process $X_{t}$ is as
where $D_{i,j}(x,t)=\frac{1}{2}\sigma(X_{t})\sigma(X_{t})^{T}$. When the Fokker–Planck equation is equal to zero, it means that the diffusion process has reached a stationary state and $p(x,t)$ is equal to the stationary distribution $\pi(X)$. Then, simulating this diffusion process is equivalent to sampling from the stationary distribution.
The proof of Proposition 3.1 and Proposition 3.2 will be concise based on the general framework for dynamics-based sampling methods proposed by Ma et al. (2015). They proved that if $\mu(X_{t})$ and $\sigma(X_{t})$ are restricted to the following form
@@ Appendix B Theoretical Proof
where $G(\cdot)$ is the potential energy function, $\mathbf{D}(\cdot)$ is a positive semidefinite matrix, and $\mathbf{Q}(\cdot)$ is a skew-symmetric curl matrix, then the Fokker–Planck equation can be transformed into a more compact form (Yin & Ao, 2006; Shi et al., 2012):
@@
Then, it is easy to verify that the stationary distribution of the diffusion process is exactly as $\pi(X)\propto\exp(-G(X))$.
@@
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ $=[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f]\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t}$ (18)
@@
where $\nabla_{f}U(f)=-\nabla_{f}\log p(\mathcal{D}|f)-\nabla_{f}\log p(f)$, $\mathbf{D}(f)=(\nabla_{\mathbf{w}}f)^{2}$, $\mathbf{Q}(f)=\mathbf{0}$, $\Gamma(f)=H_{\mathbf{w}}f$. Therefore, the Fokker-Planck equation of the probability density $p(f,t)$ for $f_{t}(\cdot;\mathbf{w})$ then can be derived as
@@
from which the stationary distribution of $p(f,t)$ can be verified as $\exp(-U(f))=p(f|\mathcal{D})$, that is, the target posterior distribution over functions.
@@
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&0\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}U(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t$ (20)
@@
where $\mathbf{Q}(f,g)$ is s a skew-symmetric curl matrix, $\mathbf{D}(f,g)=\mathbf{0}$. Therefore, the stationary distributions of the functional Hamiltonian dynamics defined in Equation 10 is $\pi(f,g)\propto\exp(H(f,g))$, and simply (marginalize) discard the auxiliary variables can obtain the target true posterior $p(f|\mathcal{D})$.
@@
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}\tilde{U}(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t+\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g)^{T}\end{array}\right]\mathrm{d}t,$ (21)
where $\mathbf{D}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$ and $\mathbf{Q}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ is the same as that in full-batch functional Hamiltonian dynamics. Then, the (marginal) stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$.
@@
Appendix C Pseudocode for Functional SGMCMC.
@@
Appendix D Ablation Study
D.1 The effects of the sample size for naive SGMCMC and functional SGMCMC

### v2

@@ 5.2 UCI Regression
In this experiment, we evaluate our methods on multivariate regression tasks across seven benchmark UCI datasets: Yacht, Boston, Concrete, Energy, Wine, Kin8nm and Protein. We use a two-hidden-layer fully connected neural network. For all sampling methods, we run 500 iterations for the burn-in stage and collect 15 samples. Performance is measured via average root mean square error (RMSE) and the test negative log-likelihood (NLL). For each dataset, we construct 10 random 90-to-10 train-test splits and report the mean and standard deviation of RMSE and NLL over these splits in Table 2 and Table 3, respectively. fSGHMC outperforms all baselines for RMSE on all datasets, and fSGLD also achieves better accuracy values than naive SGMCMC methods and functional variational inference, demonstrating superior performance of our functional SGMCMC in predictive accuracy. Moreover, fSGHMC and fSGLD obtain better NLL results on most of the datasets (5/7), showing that our methods are also very competitive in uncertainty quantification.
Table 2: The table shows the results of average RMSE for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity. Bold indicates statistically significant best results ($p<0.01$ with t-test).
@@
Table 3: The table shows the results of average NLL for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity. Bold indicates statistically significant best results ($p<0.01$ with t-test).
@@ 5.3 Imagine Classification and OOD Detection
In this section, we investigate the scalability of our methods on image classification tasks with high-dimensional inputs. We evaluate the in-distribution predictive performance and out-of-distribution (OOD) detection ability on MNIST (LeCun et al.,, 2010), FashionMNIST(Xiao et al.,, 2017) and CIFAR-10(Krizhevsky et al.,, 2009). We use ResNet-18 architecture (He et al.,, 2016) for all methods on these three datasets, and the batch size is 125. For all sampling methods, we run for 60 burn-in iterations and collect 10 samples on the MNIST dataset, 100 burn-in iterations and collect 10 samples on FashionMNIST, 80 burn-in iterations and draw 10 samples on CIFAR10. We randomly sample 40 points from training data in every iteration as the measurement points for the gradient estimation of log functional prior distribution in fSGLD and fSGHMC. We report the test classification errors for predictive performance and the area under the curve (AUC) of OOD detection pairs FashionMNIST/MNIST, MNIST/FashionMNIST and CIFAR10/SVNH based on predictive entropies in Table 4. Our fSGLD and fSGHMC outperform all other methods for both predictive accuracy and OOD detection.
@@ 5.4 Contextual Bandits
In many downstream tasks, such as online sequential decision-making scenarios in contextual bandits problems, reliable uncertainty estimation is important to guide the exploration-exploitation process. In such problems, an agent interacts with an unknown environment repeatedly and chooses to take an optimal action in each round of interaction. Thompson sampling (Thompson,, 1933; Russo and Van Roy,, 2016) is a widely used algorithm for exploration strategies in contextual bandits, it (i) first samples a current posterior to obtain a model configuration; (ii) then adaptively chooses an optimal action based on the current context under the sampled model configuration and observes the corresponding reward; (iii) updates the posterior based on the context, action and reward tuples in this interaction round.
This section evaluates the ability to guide exploration on the UCI Mushroom dataset with 8124 instances. In each instance, the mushroom has 22 features and is labeled as edible or poisonous. An agent can observe mushroom features as the context in each interaction and choose to eat or reject a mushroom. We compare our fSGLD with several baselines, including naive SGLD, functional variational FBNN and IFBNN, parametric-space variational method Bayes By Backprop (BBB) (Blundell et al.,, 2015). We follow the basic settings by Blundell et al., (2015) and consider three different reward patterns: if the agent eats an edible mushroom, it receives a reward of 5 and a reward of 0 if the agent chooses to reject. On the other hand, if the agent eats a poisonous mushroom, it receives a reward of -35 with probabilities of 0.4, 0.5, and 0.6, respectively, for three different patterns. If the agent chooses to reject a poisonous mushroom, it receives a reward of 0. Suppose an oracle always chooses to eat an edible mushroom and reject the poisonous mushroom. The cumulative regrets with respect to the reward achieved by the oracle measure the exploration-exploitation ability of an agent. We run 10000 iterations for all methods and the minibatch size is 64. Figure 2 shows the cumulative regrets for all methods in three reward patterns. Our fSGLD consistently outperforms other inference methods, indicating that our functional SGMCMC can provide reliable uncertainty estimation in decision-making tasks.
@@ 6 Discussion and Conclusions
In this paper, we lift the Itô diffusion of network parameters in BNNs onto function space and propose novel functional MCMC schemes for BNNs. Specifically, we design functional Langevin dynamics and functional Hamiltonian dynamics for the function mapping defined by BNNs and derive corresponding tractable sampling methods in parameter space, which can incorporate informative functional prior into posterior inference. Moreover, we proved that the stationary measure of our functional dynamics is the posterior over functions. Empirically, we demonstrate that our functional MCMC schemes effectively leverage functional priors to yield superior predictive performance and principled uncertainty quantification on several benchmark tasks.
Appendix A Notation Table
Table 5 is the notation table to demonstrate the notation used in this paper.
Table 5: Notation table
Notation Meanings
$\mathcal{D}=\left\{\mathbf{X}_{\mathcal{D}},\mathbf{Y}_{\mathcal{D}}\right\}$ Training dataset
$\mathcal{X}\subseteq\mathbb{R}^{p}$ ($p$-dimensional) input space
$\mathcal{Y}\subseteq\mathbb{R}^{c}$ ($c$-dimensional) output space
$(\Omega,\mathcal{A},P)$ Probability space on $\mathbb{R}^{k}$
$\mathbb{H}$ Infinite-dimensional function space (Polish space)
$\mathcal{B}(\mathbb{H})$ Borel $\sigma$-algebra on $\mathbb{H}$
$\mathcal{P}(\mathbb{H})$ The space of Borel probability measures on $\mathcal{B}(\mathbb{H})$.
$X$ Stochastic process
$\mathbf{X}_{\mathcal{M}}$ Finite measurement points from input space
$\mathbf{w}\in\mathbb{R}^{k}$ Random network parameters
$\mathbf{z}\in\mathbb{R}^{k}$ Momentum variable
$\epsilon$ Discretization step size
$\eta$ Standard Gaussian noise
$f(\cdot;\mathbf{w})$ Random function mapping defined by a BNN parameterized by $\mathbf{w}$
$p_{0}(\mathbf{w})$ Prior distribution over network parameters
$p(\mathbf{w}|\mathcal{D})$ Posterior distribution over network parameters
$p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$ Likelihood function evaluated on the training data
$p(\mathbf{z})$ Auxiliary probability distribution
$P_{0}$ Functional Prior measure
$\operatorname{supp}(P_{0})$ Topological support of $P_{0}$
$E$ A non-empty subset of $\operatorname{supp}(P_{0})$
$P_{f|\mathcal{D}}$ Posterior measure over functions
$\Phi(f)$ Negative log-likelihood as $-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$
$p(g)$ Functional auxiliary probability measure
$U(\mathbf{w})$ Potential energy function in parameter-space Langevin dynamics
$H(\mathbf{w},\mathbf{z})$ Potential energy function in parameter-space Hamiltonian dynamics
$I_{0}(f)$ Onsager–Machlup (OM) functional for $P_{0}$
$U(f)$ Potential energy functional in functional Langevin dynamics
$H(f,g)$ Potential energy functional in functional Hamiltonian dynamics
$\mu(\cdot)$ Drift term in the SDE of Itô diffusion
$\sigma(\cdot)$ Diffusion coefficient in the SDE of Itô diffusion
$B$ Standard Wiener process (Brownian motion)
$\pi(\cdot)$ Stationary distribution (measure)
Appendix B Further Background
Itô diffusion
Dynamics-based MCMC methods are rooted in the general framework of Itô diffusion (Øksendal,, 2003), which is commonly used to model the evolution of particles in a system. Given a stochastic process $X:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}$ defined on a probability space $(\Omega,\Sigma,P)$, an Itô diffusion in n-dimensional Euclidean space, driven by the standard Wiener process (Brownian motion), satisfies the following specific type of stochastic differential equation (SDE):
$\mathrm{d}X_{t}=\mu(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ (11)
where $X_{t}$ represents the state of the stochastic process at time $t$, $\mu(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a vector field describing the deterministic drift term, and $\sigma(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ is a matrix field denotes the diffusion coefficient for $X_{t}$. Both $\mu(\cdot)$ and $\sigma(\cdot)$ are assumed to satisfy the standard Lipschitz continuity condition (Ghosh,, 2010). The term $B\in\mathbb{R}^{n}$ refers to an n-dimensional Wiener process (Brownian motion), with $\mathrm{d}B_{t}$ being the increment of a Wiener process, distributed as $\mathrm{d}B_{t}\sim\mathcal{N}(0,\mathrm{d}t\cdot\mathbf{I}_{n})$. The probability density function $p(x,t)$ of $X_{t}$ is governed by the Fokker-Planck (FP) equation (Risken,, 1996):
where $D_{i,j}(x,t)=\frac{1}{2}\sigma(X_{t})\sigma(X_{t})^{T}$. This equation serves as the foundation for deriving dynamics-based MCMC sampling methods. The diffusion process reaches a stationary state when the FP equation equals zero, then simulating this diffusion process is equivalent to sampling from the stationary distribution, denoted by $\pi(x)$. By carefully designing $\mu(\cdot)$ and $\sigma(\cdot)$, one can sample from the target distribution as the stationary distribution. For example, for the commonly used Langevin dynamics, where $\mu(X)=-\nabla U(X)$, $\sigma(X)=\sqrt{2}\cdot\mathbf{I}_{n}$, and $U(\cdot)$ is a potential energy function. The stationary distribution for $p(x,t)$ is given by the Boltzmann distribution $\pi(x)\propto\exp(-U(X))$. Thus, when sampling the posterior $p(\mathbf{w}|\mathcal{D})$ in BNNs, the potential energy function is given by $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})-\log p(\mathbf{w})$, and $\sigma(\mathbf{w})=\sqrt{2}\cdot\mathbf{I}_{k}$, ensuring that the target posterior as the stationary distribution.
Itô Lemma
@@ Appendix B Further Background
Itô Lemma (Itô,, 1951) is a fundamental result in stochastic calculus to find the differential of a function of a stochastic process, which serves as the stochastic calculus counterpart of the chain rule. For an Itô diffusion, let $f(X_{t})$ be an arbitrary twice differentiable scalar function of real variables $X_{t}$, then the differential of $f(X_{t})$ can be derived from the Taylor series expansion of the function as
@@
which implies that $f(X_{t})$ is itself an Itô diffusion (Brzeźniak et al.,, 2008).
Onsager–Machlup functional
(see e.g., Lambley, (2023), Definition 2.4.) Suppose $\mathbb{H}$ is a separable metric space and $P_{0}\in\mathcal{P}(\mathbb{H})$. Assume that $\varnothing\neq E\subseteq\operatorname{supp}(P_{0})$, then a functional $I_{0}:E\rightarrow\mathbb{R}$ is an Onsager–Machlup (OM) functional for $P_{0}$ if for all $f,f^{\prime}\in E$,
$\lim_{r\rightarrow 0}\frac{P_{0}(B_{r}(f)}{P_{0}(B_{r}(f^{\prime})}=\exp(I_{0}(f^{\prime})-I_{0}(f)),$ (14)
where $B_{r}(\cdot)$ denotes the closed ball of radius $r$ in $\mathbb{H}$ and $\operatorname{supp}(P_{0}):=\{f\in\mathbb{H}\ |\ P_{0}(B_{r}(f))>0\ \text{for all}\ r>0\}$. Lambley, (2023) (in Proposition 2.6.) proved that posterior measure $P_{f|\mathcal{D}}\in\mathcal{P}(\mathbb{H})$ defined as $P_{f|\mathcal{D}}(\mathrm{d}f)\propto\exp(-\Phi(f))P_{0}(\mathrm{d}f)$ has OM functional $I^{y}:E\rightarrow\mathbb{R}$ given by $I^{y}(f)=I_{0}(f)+\Phi(f)$, where $\Phi(f)$ is the continuous potential, in essence the negative log-likelihood as $-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$. OM functional can be interpreted heuristically as the negative logarithm of the Lebesgue density, but this interpretation cannot be taken literally in infinite-dimensional spaces, as there is no Lebesgue measure in such settings. For instance, the OM functional for a Gaussian measure on an infinite-dimensional Banach space is typically defined only on a specific subspace, known as the Cameron-Martin space, e.g., for a centred Gaussian measure, the OM functional is defined as $I:E\rightarrow\mathbb{R},I(f)=\frac{1}{2}||f||^{2}_{E}$. As a result, the OM functional does not need to be defined over the entire space $\mathbb{H}$.
Appendix C Theoretical Proof
The proof of Proposition 3.1 and Proposition 3.2 will be concise based on the general framework for dynamics-based sampling methods proposed by Ma et al., (2015). They proved in Theorem 1 that if $\mu(X_{t})$ and $\sigma(X_{t})$ are restricted to the following form
@@ Appendix C Theoretical Proof
where $G(\cdot)$ is the potential energy functional, $\mathbf{D}(\cdot)$ is a positive semidefinite matrix, and $\mathbf{Q}(\cdot)$ is a skew-symmetric curl matrix, then the Fokker–Planck equation can be transformed into a more compact form (Yin and Ao,, 2006; Shi et al.,, 2012) given by
@@
Then, it is straightforward to verify that the stationary distribution of the diffusion process is exactly as $\pi(X)\propto\exp(-G(X))$.
@@
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ $=[-(\nabla_{\mathbf{w}}f_{t})^{T}(\nabla_{\mathbf{w}}f_{t})\nabla_{f}U(f_{t})+H_{\mathbf{w}}f]\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f_{t})^{T}\mathrm{d}B_{t}$ (17)
@@
where $U(f)=-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))+I_{0}(f)$ is an OM functional for $P_{f|\mathcal{D}}$, $\mathbf{D}(f)=(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)$, $\mathbf{Q}(f)=\mathbf{0}$, $\Gamma(f)=H_{\mathbf{w}}f$. Therefore, the Fokker-Planck equation of the probability measure $p(f,t)$ for $f_{t}(\cdot;\mathbf{w})$ then can be derived as
@@
from which the stationary probability measure of $p(f,t)$ can be verified as $\exp(-U(f))=P_{f|\mathcal{D}}$, that is, the target posterior measure over functions.
@@
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g_{t}\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f_{t}&0\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}U(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t$ (19)
@@
where $\mathbf{Q}(f,g)$ is s a skew-symmetric curl matrix, $\mathbf{D}(f,g)=\mathbf{0}$. Therefore, the stationary measures of the functional Hamiltonian dynamics defined in Equation 7 is $\pi(f,g)\propto\exp(H(f,g))$, and simply (marginalize) discard the auxiliary functions can obtain the target true posterior $P_{f|\mathcal{D}}$.
@@
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g_{t}\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f_{t}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g_{t}\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}\tilde{U}(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t+\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g_{t})^{T}\end{array}\right]\mathrm{d}t,$ (20)
where $\mathbf{D}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$ and $\mathbf{Q}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ is the same as that in full-batch functional Hamiltonian dynamics. Then, the (marginal) stationary probability measure of such functional stochastic gradient Hamiltonian is still the target functional posterior $P_{f|\mathcal{D}}$.
@@
Appendix D Pseudocode for Functional SGMCMC.
@@
Appendix E Experimental Setting
Extrapolation on Synthetic Data
In this experiment, we use $2\times 100$ fully connected tanh neural networks for all models. The functional GP prior with the RBF kernel is pre-trained on the 20 training points for 100 epochs for all functional methods. we use a decreasing step-size schedule for $\epsilon_{t}$ with 1e-3 as the initial value for all sampling methods. The specific decreasing schedule is as every 5000 iterations, the decay factor is 0.9.
Multivariate regression on UCI datasets
In this experiment, we use two-hidden-layer fully connected neural networks and each layer with 10 hidden units for all models. The functional GP prior with the RBF kernel is pre-trained for 100 epochs. We run 500 iterations for the burn-in stage and collect 15 samples in the following 1500 iterations for all sampling methods. The initial $\epsilon_{t}$ is 1e-3, and the decay factor is 0.9. Functional variational inference methods are trained for 2000 epochs for fair comparison.
Imagine Classification and OOD Detection
In this experiment, we use ResNet-18 architecture for all methods. The functional prior is a Dirichlet-based GP designed for classification tasks and is pre-trained for 500 epochs. The initial $\epsilon_{t}$ is 1e-2, the decay period is 20, and the decay factor is 0.99. For all sampling methods, we run for 60 burn-in iterations and collect 10 samples in the following 20 iterations on the MNIST dataset, 100 burn-in iterations and collect 10 samples in the following 100 iterations on Fashion-MNIST, 80 burn-in iterations and draw 10 samples in the following 20 iterations on CIFAR10. Functional variational inference methods are trained for 80, 200, and 100 epochs on MNIST, FashionMNIST and CIFAR-10, respectively.
Contextual Bandits
In this experiment, we use fully connected neural networks with input-100-100-output architecture for all methods. The GP prior with RBF kernel is pre-trained on 1000 randomly sampled points from training data for 100 epochs. The initial $\epsilon_{t}$ is 1e-2, and the decay factor is 0.9. ALL models are trained using the last 4096 input-output tuples in the training buffer with a batch size of 64 and training frequency of 64 for each iteration.
Appendix F Analysis of Mixing Time
The trajectories of our fSGLD and the naive parameter-space SGLD for both the 1-D extrapolation and UCI regression (Yacht) experiments are shown in Figure 3. For the extrapolation example, we use a two-hidden-layer fully connected neural network with 100 hidden units in each layer (resulting in a 10401-dimensional parameter space). For the UCI regression task, we use a 2×10 fully connected network (141-dimensional parameter space). As illustrated in Figure 3, our fSGLD converges rapidly to the stationary measure—within 2000 iterations for the extrapolation example and 500 iterations for the UCI regression task. This convergence rate is comparable to that of the parameter-space SGLD, highlighting the rapid mixing rate of our method.
(a) fSGLD-toy
(b) fSGLD-uci
(c) SGLD-toy
(d) SGLD-uci
Figure 3: Log-posterior probability versus the number of iterations.
Appendix G Computational Complexity
The runtime comparisons of a single iteration for all methods on the 1-D extrapolation, UCI regression (yacht) and image classification (MNIST) are provided in Table 7. For our fSGLD, fSGHMC, functional variational IFBNN and FBNN methods, the functional GP prior was pre-trained for 100, 100, and 500 epochs, respectively, in these three experiments, which added only an extra 1s for both the 1-D and UCI experiments, and 3s for the classification task. Compared to functional variational inference methods, our functional SGMCMC schemes exhibit significantly better computational performance, and are only slightly slower than the naive parameter-space SGMCMC. This demonstrates the computational efficiency of our method while maintaining the advantages of incorporating functional priors for improved inference.
Table 7: Runtime comparison of single iteration for all methods (in seconds).
Model 1-D extrapolation UCI Image classification
SGLD 0.0016 0.0025 4
fSGLD 0.0044 0.0045 5.83
SGHMC 0.0023 0.02 10.95
fSGHMC 0.0051 0.044 13.64
IFBNN 0.0094 0.0075 6.81
FBNN 0.129 0.114 23.25
Appendix H Ablation Study
H.1 The effects of the sample size for naive SGMCMC and functional SGMCMC